Need to solve the following
8y^2=14y-5 (hint set equal to 0 first)

Answer by stanbon(57224)

to solve the following
8y^2=14y-5
8y^2 - 14y + 5 = 0
---
Think of two numbers whose product is 5*8=40
and whose sum is -14.
---------------
They are -10 and -4
---------------
Rewrite the problem:
8y^2 - 10y - 4y + 5 = 0
---
Factor the 1st two and the last two terms seperately:
2y(4y-5) - (4y-5) = 0
Factor again:
(4y-5)(2y-1) = 0
y = 5/4 or y = 1/2
============================
Cheers,
Stan H.